Find the sum of squares of all numbers x such that both expressions x^2+5x and x+(1/x) are integer numbers.
1
Expert's answer
2020-12-01T01:58:32-0500

"let \ x+\dfrac{1}{x} = k , k- integer\ number\\
\dfrac{x^2+1}{x} = k\\
x^2-kx+1 = 0\\
x = \dfrac{k \pm \sqrt{k^2-4}}{2}\\
let \ x^2+5x = n\\
x^2+5x-n = 0\\
x = \dfrac{-5 \pm \sqrt{25-4n}}{2}\\
\dfrac{k \pm \sqrt{k^2-4}}{2} = \dfrac{-5 \pm \sqrt{25-4n}}{2}\\
k \pm \sqrt{k^2-4} = -5 \pm \sqrt{25-4n}\\
\text{which equations have 3 integer solutions} \\
k = -5 , n = 1\\
k = -2, n = 4\\
k = 2, n = -6\\
x = 1, x=-1 , x = \dfrac{-5 \pm \sqrt{21}}{2}\\
S= 1^2+(-1)^2 + (\dfrac{-5 + \sqrt{21}}{2})^2 +(\dfrac{-5 - \sqrt{21}}{2})^2 = \\
= 1+1 +(\dfrac{25 + -10\sqrt{21} +21 +25+10\sqrt{21}+21}{4}) = 25
\\answer: 25"